A non-relativistic free electron has kinetic energy K. If its wavelength λ triples, it is kinetic energy is:
1
Expert's answer
2018-04-06T05:16:11-0400
The de Broglie wavelength of electron is
λ=h/p=h/mv=h/√2mK, (1)
since K=〖mv〗^2/2.
From (1) we obtain K=h^2/(2mλ^2 ).
So if the wavelength λ triples, kinetic energy K decreases 9 times.